OpenStudy (anonymous):

Choose the correct degrees and minutes for: 63.6°

jimthompson5910 (jim_thompson5910):

you take the decimal portion 0.6 and multiply it by 60 (since there are 60 min in a degree) to get 0.6*60 = 36 so 36 minutes = 0.6 degrees

jimthompson5910 (jim_thompson5910):

there are no decimal portions to the result we got of 36, so there are no seconds (or you could say there are 0 seconds)

jimthompson5910 (jim_thompson5910):

this means 63.6 degrees =63 degrees, 36 minutes, 0 seconds or simply 63.6 degrees =63 degrees, 36 minutes

OpenStudy (anonymous):

:D thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

what about 18.27?

jimthompson5910 (jim_thompson5910):

0.27*60 = ???

OpenStudy (anonymous):

16.2

jimthompson5910 (jim_thompson5910):

then you multiply the decimal portion of that result by 60 0.2*60 = 12

jimthompson5910 (jim_thompson5910):

this means 18.27 degrees = 18 degrees, 16 minutes, 12 seconds

OpenStudy (anonymous):

ohh okay, that makes sense

jimthompson5910 (jim_thompson5910):

that's great, glad it does